please answer my question...i really need this ASAP...
"There are 2 numbers which has a sum of 52. If the larger number is divided to the smaller number and their quotient is 9 and their remainder is 2...what is the number???

Answer by mananth(16946)

larger number =x
smaller number = y
x+y =52................(1)
(x-2)/y = 9
x-2 = 9y
x-9y=2.................(2)
subtract (2) from (1)
x-x+y-(-9y)=52-2
y+9y=50
10y=50
y= 5
x+y=52
y=5
therefore x= 47
47 & 5